# Grassroots Consensus

### v2

@@ 2.3 The Protocol
In the Grassroots Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$. The protocol is event-driven as described in Algorithm 1.
Overview. Each three consecutive rounds are grouped into a wave, which aims to finalize one block of its first round. Similarly to Morpheus [28], the protocol has two modes of operation, high-throughput and low-throughput, but here they are implicit— depending on whether the preceding wave is quiescent. In high-throughput mode, in the good-case scenario, only the leader sends a first round block. If the leader is faulty or does not communicate with other agents in a timely manner (bad-case scenario), other agents time-out on the leader and send blocks of their own. Agents wait either for a first-round leader block or for a supermajority of blocks in a round before producing a block in the next round.
Low-throughput is manifested by a quiescent wave, which finalizes its first-round spontaneous leader using two rounds of empty blocks. Following a quiescent wave, any agent may initiate a new wave by issuing a non-empty block. If two such blocks collide the protocol enters high throughput mode until, if ever, it quiesces again.
Initialize: $D\leftarrow\phi$, $B\leftarrow\phi$, $\emph{payload}\leftarrow\bot$.
Let $r$ always be the maximal advanced round in $B$, which is initially 0.
Receive: Upon receipt of $q$-block $b$, if $b$ a nack-block then judiciously send $[b]$ to $q$,
else if $b$ a nudge-block such that $p=\emph{leader}(d(b))$ and $[b]\not\subseteq B\cup\{b\}$ then send once a nack-block for $b$ to $q$,
else if $b$ is valid then add $b$ to $D$.
Accept & Nack: If $[b]\subseteq B\cup\{b\}$ for some $b\in D$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$,
else if a $q$-block $b^{\prime}$ was added to $D$ more than $\Delta$ time ago then send once a nack-block for $b^{\prime}$ to $q$.
@@
Issue: Issue once a new block if $r+1$ is a second or third round or a first round that follows a wave that is: 1. quiescent (or $r=0$) and $\emph{payload}\neq\bot$, or 2. non-quiescent and (i) $p=\emph{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$. Nudge: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once to $\emph{leader}_{P}(r+1)$ a nudge-block for round $r+1$. Output: If $B$ contains a final block $b$ of depth greater than the previous final block (if any) then output every block in $\tau(b)$ unless it has already been output.
@@
issue a new block means create a block $b$ with payload and pointers to the tips of $B_{r}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
@@
Algorithm 1 Grassroots consensus protocol, code for agent $p\in P$, given $(P,\sigma,\Delta)$.
@@
$2\Delta$ for nudging the ensuing round’s leader. In the good case whence all agents are correct after GST, whenever an agent receives a block, all other agents receive it within $\Delta$ time. Thus, once a third round is advanced at the blocklace of a correct non-leader agent $p$, $p$ expects to receive a leader block of the ensuing (first) round within $2\Delta$. If this does not happen, this could be because (i) it is before GST, (ii) the leader is incorrect, or (iii) the leader has not seen yet that the previous round is advanced. This last third case can happen after GST even if the leader is correct, due to Byzantine partial dissemination of third-round blocks. Since $p$ knows that the previous third round is advanced, it sends a nudge-block to the leader, which will cause the leader to send a nack-block requesting the blocks it is missing. This $2\Delta$ delay is an optimization to avoid spurious nudges in the good case; eliminating it does not hamper safety or liveness.
@@
$9\Delta$ for an unresponsive leader. Following a non-quiescent wave after GST, if the leader $p_{\ell}$ is correct, t all correct agents should wait for $p_{\ell}$’s first round block before proceeding to the second round in order to allow that block to be finalized. But because $p_{\ell}$ may be faulty, they need to eventually time-out and stop waiting. To determine the time-out, we consider the worst-case latencies as follows: Assume a third round $r$ is advanced at $p$ at time $t$. If $p$ does not receive a leader block by $t+2\Delta$, it nudges $p_{\ell}$, which issues a nack-block, upon receipt of which $p$ sends $p_{\ell}$ any missing blocks. This exchange takes at most $3\Delta$, so by time $t+5\Delta$, $p_{\ell}$ issues a leader block $b$ for round $r+1$, which is received by $p$ by time $t+6\Delta$. Upon receipt of $b$, $p$ might be missing some blocks in $[b]$ and wait $\Delta$ before sending a nack and getting them from $q$ by time $t+9\Delta$. If $p$ does not accept any leader-block by $t+9\Delta$, it concludes that either $p_{\ell}$ is faulty or GST has not arrived yet, and sends a first-round block of its own in order to allow the protocol to proceed to the next round.
@@
Protocol walk-through. Incoming blocks are inserted to $D$; if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $B$.
The variable $r$ always holds the highest advanced round in $B$. If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block. If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode. During low throughput (initially and following a quiescent wave), $p$ issues a block only if it has payload to send. During high-throughput, $p$ issues a block in case (i) it is the leader, or (ii) it times-out on the leader.
@@
An agent $p$ always issues blocks that follow the latest advanced round $p$ saw. This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced. This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own.
After GST, whenever there is a wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block. Therefore, no agent can produce a valid second round block that does not endorse the leader block. So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized.
@@
Ordering with $\tau$. Once the protocol finalizes a block $b$, it orders the predecessors of $b$ that have not been ordered yet using the function $\tau$. The definition of $\tau$ is similar to Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks.
The output of $\tau(b)$ is a sequence that includes all blocks observed by $b$, with equivocating blocks excluded. In Section 3 we prove that $\tau$ is monotonic and safe. The monotonicity of $\tau$ ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor. Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any. The safety of $\tau$ means that blocks are ordered the same way at all correct agents.
In anticipation of the protocol being constitutional, we already assume the notion of the prevailing constitution. For now, assume $(P,\sigma,\Delta)$ is fixed arbitrarily for all blocks. In Section 6, we define the constitution prevailing in a block, and have each call (initial and recursive) to $\tau$ with a block $b$ employ the parameters of the constitution prevailing in $b$.
@@
We assume a fixed topological sort function $\textit{xsort}(b,B)$ (exclude and sort) that takes a block $b$ and a blocklace $B$, and returns a sequence that preserves the ‘observes’ partial order and includes all the blocks in $B$ that are approved by $b$. The function $\tau$ takes a block as input and returns a sequence of blocks in $[b]$ as output, as follows:
$\tau(b):=\begin{cases}\tau(b^{\prime})\cdot\textit{xsort}(b,[b]\setminus[b^{\prime}])\text{\ \ if $b^{\prime}$ is the maximal-depth block ratified in $[b]$}\\ \ \ \ \ \ \ \ \ \ \textit{according to the constitution prevailing in $b$}\\ \textit{xsort}(b,[b])\text{\ \ \ \ \ if no such $b^{\prime}$ exists }\\ \end{cases}$
@@
Latency and complexity. Regarding latency, we note that after GST and in the good case a (formal or spontaneous) leader block is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
@@
Regarding communication complexity, during low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all. In Morpheus [28], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$. A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead. Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block. A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, should carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block. Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity. To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.k
@@
3 Safety and Liveness of Grassroots Consensus
Here we prove the safety and liveness of the Grassroots Consensus protocol.
We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
Definition 3.1 (Safety and Liveness of an Ordering Consensus Protocol).
An ordering consensus protocol is:
Safe if output sequences of correct agents are always consistent.
Live if every non-empty block issued by a correct agent is eventually included in the output of every correct agent.
Next we prove safety and liveness of the basic Grassroots Consensus protocol.
@@ 3.1 Safety
Grassroots Consensus is safe.
To prove the theorem, we follow a condensed and simplified proof of the safety of Cordial Miners.
First, observe that Grassroots Consensus produces only $f$-safe blocks; if a block created by the protocol is not $f$-safe it means that its closure includes equivocations by more than $f$ agents, a contradiction. For the following proofs we generalize $\tau$ to blocklaces, so that $\tau(B)$ is $\tau(b)$ if $b$ is the most-recent final block in $B$ if there is one, else it is the empty sequence $\Lambda$.
@@
A blocklace $B$ is $\tau$-safe if every final block in $B$ is ratified by every higher-depth block ratified by some block in $B$.
@@
An f-safe blocklace is $\tau$-safe.
Proof 3.5.
Let $B$ be an $f$-safe blocklace, let $b$ be a block of wave $k$ final in $B$ and $b^{\prime}$ a block of a later wave $k^{\prime}>k$ ratified by some block in $B$, which means that $b^{\prime}$ is endorsed by a $\sigma$-supermajority in $B$. We have to show that $b$ is ratified by $b^{\prime}$.
@@
Then the blue supermajority and the finalizing $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection. Namely, there is a correct agent $p$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $k^{\prime}$ happens to be $k+1$), which in turn must observe a supermajority of blocks that endorse $b$. Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$.
@@
$\tau$ is monotonic wrt $\supset$ over $f$-safe blocklaces.
Proof 3.7.
We have to show that if $B\subset B^{\prime}$ for $f$-safe blocklaces $B,B^{\prime}$ then $\tau(B)$ is a prefix of $\tau(B^{\prime})$. By Observation 2.2, all the blocks final in $B$ are also final in $B^{\prime}$. By Observation 2.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their round number. If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially. Assume $b\neq b^{\prime}$ are the most-recent final blocks in $B,B^{\prime}$, respectively, so necessarily $b^{\prime}$ is more recent than $b$. Let $b=b_{1},\ldots,b_{k}=b^{\prime}$, $k\geq 2$, be the sequence of all ratified blocks in $B^{\prime}$ between and including $b$ and $b^{\prime}$. By Proposition 3.4 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
@@
If the union $B\cup B^{\prime}$ of two blocklaces is $f$-safe then $\tau(B)$ and $\tau(B^{\prime})$ are consistent.
Proof 3.8.
By monotonicity of $\tau$ over safe blocklaces, $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
Proof 3.9 (Proof of Theorem 3.2).
Since in a run of Grassroots Consensus there are at most $f$ faulty agents in total, the union of the blocklaces of two correct agents is always $f$-safe. Hence, by Observation 3.1, the outputs of any two correct agents running Grassroots Consensus are always consistent.

### v4

@@ 3.3 Single-Epoch Constitutional Consensus Pseudocode
In the Constitutional Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$. An epoch identifier $e$ points to the epoch’s constitutional genesis block. We use $B^{e}$ to denote the blocklace of epoch $e$ in $B$, which is the subset of $B$ that observes $e$. The protocol is event-driven. Algorithm 1 describes a single-epoch Constitutional Consensus, in which the prevailing constitution is fixed throughout the run.
Initialisation. The first epoch is initialised with $D\leftarrow\phi$, $B\leftarrow\{b\}$, $e\leftarrow b$, where $b$ is a constitutional genesis block and $\emph{payload}\leftarrow\bot$. The initialisation of subsequent epochs is discussed in Algorithm 2.
Let $B:=B^{e}$ herein and $r$ be the maximal advanced round of the current $B^{e}$. Receive: Upon receipt of a $q$-block $b$, $q\in P$
if $b$ is a nack-block for a $p$-block then judiciously send $[b]$ to $q$,
else if $b$ an inform-block and $b]\not\subseteq B\cup\{b\}$
——- then send once a nack-block for $b$ to $q$,
else if $b$ is a valid ordinary block then $D\leftarrow D\cup\{b\}$.
Accept or Nack: For every $q$-block $b\in D$, $q\in P$,
if $b]\subseteq B\cup\{b\}$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$,
else if a $b$ was added to $D$ more than $\Delta$ time ago
——- then send once a nack-block for $b^{\prime}$ to $q$.
@@
Issue: Issue once a round $r+1$ block if $r+1$ is a second or third round or a first round that follows a wave that is: 1. quiescent and $\emph{payload}\neq\bot$, or 2. non-quiescent and (i) $p=\textit{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$. Issue Backlog: If $\emph{payload}\neq\bot$ and no round $r$ or round $r+1$ block has been issued by $p$, then issue once a round $r$ block. Inform Leader: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once an inform-block for round $r+1$ to $\textit{leader}_{P}(r+1)$. Output: If $B$ contains a final block $b$ of depth greater than the previous final block then output every block in $\tau(b)$ (Definition 3.8) unless it has already been output.
@@
issue a round $k$ block means create a block $b$ with payload and pointers to the $P$-blocks that are tips of $B_{k-1}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
@@
Algorithm 1 Constitutional Consensus for a single epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$. Code for agent $p\in P$.
@@
$2\Delta$ for informing the ensuing round’s leader. In the good case whence all agents are correct after GST, whenever an agent receives a block, all other agents receive it within $\Delta$ time. Thus, once a third round is advanced at the blocklace of a correct non-leader agent $p$, $p$ expects to receive a leader block of the ensuing (first) round within $2\Delta$. If this does not happen, this could be because (i) it is before GST, (ii) the leader is incorrect, or (iii) the leader has not seen yet that the previous round is advanced. This last third case can happen after GST even if the leader is correct, due to Byzantine partial dissemination of third-round blocks. Since $p$ knows that the previous third round is advanced, it sends an inform-block to the leader, which will cause the leader to send a nack-block requesting the blocks it is missing. This $2\Delta$ delay is an optimization to avoid spurious informs in the good case; eliminating it does not hamper safety or liveness.
@@
$9\Delta$ for an unresponsive leader. Following a non-quiescent wave after GST, if the leader $p_{\ell}$ is correct, all correct agents should wait for $p_{\ell}$’s first round block before proceeding to the second round in order to allow that block to be finalized. But because $p_{\ell}$ may be faulty, they need to eventually time-out and stop waiting. To determine the time-out, we consider the worst-case latencies as follows: Assume a third round $r$ is advanced at $p$ at time $t$. If $p$ does not receive a leader block by $t+2\Delta$, it informs $p_{\ell}$, which issues a nack-block, upon receipt of which $p$ sends $p_{\ell}$ any missing blocks. This exchange takes at most $3\Delta$, so by time $t+5\Delta$, $p_{\ell}$ issues a leader block $b$ for round $r+1$, which is received by $p$ by time $t+6\Delta$. Upon receipt of $b$, $p$ might be missing some blocks in $[b]$ and wait $\Delta$ before sending a nack and getting them from $q$ by time $t+9\Delta$. If $p$ does not accept any leader-block by $t+9\Delta$, it concludes that either $p_{\ell}$ is faulty or GST has not arrived yet, and sends a first-round block of its own in order to allow the protocol to proceed to the next round.
Protocol walk-through.
Receive: valid incoming ordinary blocks are inserted to $D$.
@@
Accept or Nack: if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $D$.
@@
Issue: The variable $r$ always holds the highest advanced round in $B$, and an agent typically issues blocks in round $r+1$. This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced. This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own. The only exception (to sending in round $r+1$) is the Issue Backlog rule: In case $p$ has payload to send, has not sent a round $r$ block, and is not supposed to send a round $r+1$ block, $p$ issues a round $r$ block to ensure its payload is disseminated.
If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block. If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode. During low throughput (i.e., following a quiescent wave), $p$ issues a block if and only if it has payload to send. During high-throughput, $p$ issues a block in case (i) it is the formal leader, or (ii) it times-out on the leader. Here, a long timeout is used ($9\Delta$), before which we try to inform the leader that the round has advanced to give it a chance to progress.
Inform Leader: Upon timing out – with a short timeout of $2\Delta$ – on the formal leader in the first round of a high-throughput wave, $p$ sends an inform block to the formal leader in order to cause it to advance in case it is correct.
@@
Output: Once the protocol finalizes a block $b$, it orders the blocks approved by $b$ that have not been ordered yet using the function $\tau$. The definition of $\tau$ is essentially-identical to that of Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks. The output of $\tau(b)$ is a sequence that includes all blocks approved by $b$. Formally:
@@
Assume a safe constitution $(P,\sigma)$ and any fixed topological sort function $\textit{xsort}(b,B,B^{\prime})$ (exclude and sort) that takes a block $b$ and two blocklaces $B\supset B^{\prime}$, and returns a sequence that preserves the ‘observes’ partial order and includes all the non-empty blocks in $B\setminus B^{\prime}$ that are approved by $b$. The function $\tau$ takes a block as input and returns a sequence of blocks in $[b]$ as output, as follows:
[…]
@@
Proposition 3.9.
If the prevailing constitution is safe then $\tau(b)$ does not include equivocations.
Proof 3.10.
Consider an equivocation $b_{1},b_{2}\in[b]$. Consider the first call $\textit{xsort}(b^{\prime},[b^{\prime}],[b^{\prime\prime}])$ for which at least one of $b_{1}$ or $b_{2}$ is in $X:=[b^{\prime}]\setminus[b^{\prime\prime}]$, and the following two cases:
  1. 1.
    Both $b_{1},b_{2}\in X$: Then $b^{\prime}$ approves neither hence neither will be in the output of xsort and hence nor in the output of $\tau$.
  2. 2.
    Only one of them is in $X$, wlog $b_{1}\in X$, and since this is the first such call to xsort it must be that $b_{2}\in[b^{\prime\prime}]$. Then $b^{\prime}$ observes both $b_{1}$ and $b_{2}$ and hence will approve neither, whereas $b^{\prime\prime}$ observes only $b_{2}$, therefore will approve it and therefore only $b_{2}$ will occur in the output of $\tau$, if at all (assuming there are no other equivocations with $b_{2}$).
In either case, not both $b_{1}$ and $b_{2}$ will be in $\tau(b)$.
Safety and liveness. We assume that the prevailing constitution is safe and live (Definition 2.2). In Section 4 we prove that $\tau$ is monotonic. This ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor. Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any. In addition, as we shall see, during a protocol run the outputs of calls to $\tau$ by correct agents always occur in the same order, which is the basis for protocol safety.
Liveness is achieved after GST. Whenever there is a post-GST wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block. Therefore, no agent can produce a valid second round block that does not endorse the leader block. So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized. Once any agent has payload to send following a quiescent wave, the agent issues a block, and either the block is finalized in the same wave or the wave is non-quiescent and the protocol switches to high-throughput mode.
@@
Latency and complexity. Regarding latency, we note that after GST and in the good case a leader block (formal or spontaneous) is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
@@
Regarding communication complexity: During low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all. In Morpheus [23], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low-throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third-round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead. Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block. A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, would carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block. Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity. To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.
@@
4 Safety and Liveness of Single-Epoch Constitutional Consensus
Here we prove the safety and liveness of the single-epoch Constitutional Consensus protocol presented in Algorithm 1. In this case, the members of the founding constitution $C=(P,\sigma,\Delta)$ participate in the protocol for the entire run. The output sequence at every correct participant begins with $C$ and does not include additional constitutions.
@@ 4.1 Safety
The safety properties of Definition 2.4 can be shown by proving the consistency of the output sequences of correct agents, formally defined as follows: We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
{observation}
If the output sequences of every two correct agents are consistent with each other, then single-epoch consensus ensures safety.
Proof 4.1.
Validity trivially holds because every agent participates in every epoch. Assuming the output sequences of correct agents are all consistent with each other, we take $T$ to be their supremum and get Consistency.
@@
If the founding constitution is safe then single-epoch Constitutional Consensus is safe.
First, recall the assumption that in a blocklace $B$ created by the protocol with prevailing constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents. We refer to blocklaces and waves satisfying this property as constitutionally-safe wrt $(P,\sigma)$.
For the following proofs we generalize $\tau$ to blocklaces, so that $\tau(B)$ is $\tau(b)$ if $b$ is the most-recent final block in $B$ if there is one, else it is the empty sequence $\Lambda$.
@@
A blocklace $B$ is $\tau$-safe if for every wave $W\subset B$ with a block $b$ final in $W$ and any subsequent wave $W^{\prime}\subset B$ with a block $b^{\prime}$ ratified in $W^{\prime}$, $b$ is ratified in $W\cap[b^{\prime}]$.
@@
A constitutional blocklace with a safe constitution is $\tau$-safe.
Proof 4.5.
Let $B$ be a constitutional blocklace with a safe constitution $(P,\sigma)$, $W\subset B$ a wave with a block $b$ final in $W$, and $W^{\prime}\subset B$ a subsequent wave with a block $b^{\prime}$ ratified in $W^{\prime}$. We have to show that $b$ is ratified in $W\cap[b^{\prime}]$.
@@
Then the blue supermajority and the finalising $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection. Namely, there is a correct agent $p\in P$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $W$ and $W^{\prime}$ are consecutive waves), which in turn must observe a $\sigma$-supermajority of blocks that endorse $b$. Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$.
@@
$\tau$ is monotonic wrt $\supset$ over constitutionally-safe blocklaces.
Proof 4.7.
We have to show that if $B\subset B^{\prime}$ for blocklaces $B,B^{\prime}$ then $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
By Observation 3.2, all the blocks final in $B$ are also final in $B^{\prime}$. By Observation 3.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their depth. If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially. Assume $b\neq b^{\prime}$ are the most-recent final blocks in $B,B^{\prime}$, respectively, so necessarily $b^{\prime}$ is more recent than $b$. Let $b=b_{1},\ldots,b_{k}=b^{\prime}$, $k\geq 2$, be the sequence of all ratified blocks in $B^{\prime}$ between and including $b$ and $b^{\prime}$. By Proposition 4.4 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
@@
If the union $B\cup B^{\prime}$ of two blocklaces is constitutionally-safe then $\tau(B)$ and $\tau(B^{\prime})$ are consistent.
Proof 4.8.
By monotonicity of $\tau$ over constitutionally-safe blocklaces (Proposition 4.6), $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
Proof 4.9 (Proof of Theorem 4.2).
By assumption, the $P$-blocks in $B\cup B^{\prime}$ include equivocations by at most $f=(2\sigma-1)n$ agents, so this set is constitutionally-safe. Hence, by Observation 4.1, the outputs of any two correct agents running Constitutional Consensus are always consistent. By Observation 4.1, the protocol is safe.